OpenStudy (goformit100):

Two coins are tossed simultaneously 500 times, and we get Two heads : 105 times One head : 275 times No head : 120 times Find the probability of occurrence of each of these events.

OpenStudy (goformit100):

@Numb3r1

OpenStudy (goformit100):

Sir I am weak at probability , Please Help Me sir.

OpenStudy (goformit100):

@Luis_Rivera

OpenStudy (goformit100):

@Eyad Help Unblocking is done.

OpenStudy (anonymous):

The Solution : ========= Lets donate the event of getting : two heads by A1 . : One Head by A2 . : No head by A3 . So, probability of A1= 105/500= ? probability of A2 =275/500= ? probability of A3 =120/500= ? Calculate them and get the answer :)

OpenStudy (goformit100):

Thank you Sir

OpenStudy (anonymous):

Yw goformit ,Don't call me Sir plz :) Call me Eyad :P